Lenow’s Drug Stores and Hall’s Pharmaceuticals are competitors in the discount drug chain store business. The separate capital structures for Lenow and Hall are presented here.

Lenow Hall
Debt @ 8% $ 300,000 Debt @ 8% $ 600,000
Common stock, $10 par 600,000 Common stock, $10 par 300,000
Total $ 900,000 Total $ 900,000
Common shares 60,000 Common shares 30,000

a. Complete the following table given earnings before interest and taxes of $34,000, $72,000, and $89,000. Assume the tax rate is 10 percent. (Negative amounts should be indicated by parentheses or a minus sign. Round your answers to 2 decimal places.)
  


b-1. What is the EBIT/TA rate when the firm's have equal EPS?

EBIT/TA rate %

b-2. What is the cost of debt?
  

Cost of debt %


b-3. State the relationship between earnings per share and the level of EBIT.
  

EPS is unaffected by financial leverage when the pretax return on assets (EBIT/TA)


c. If the cost of debt went up to 10 percent and all other factors remained equal, what would be the break-even level for EBIT?
  

Break-even level:
Answer #1

Please rate..

Answer a.
i ii iii=i/ii
EBIT Total asset EBIT/TA Lenow EPS Hall EPS
34000 900000 3.78%                          0.15        (0.42)
72000 900000 8.00%                          0.72          0.72
89000 900000 9.89%                          0.98          1.23
Computation of EPS
Lenow Hall
i EBIT 34000 72000 89000 34000 72000 89000
ii interest 24000 24000 24000 48000 48000 48000
iii=i-ii EBT 10000 48000 65000 -14000 24000 41000
iv=iii*.9 PAT=EBT*(1-0.1) 9000 43200 58500 -12600 21600 36900
v Share out standing 60000 60000 60000 30000 30000 30000
vi=iv/v EPS              0.15          0.72                          0.98        (0.42)          0.72          1.23
Answer b.-1
at 8% EBIT/TA both the firm has equal EPS $0.71 per share
Answer b.-2
Cost of debt =8%
Answer b.-3
relationship
Pretax return is lower than cost of debt therefore Hall EPS is lower compared to Lenow
Pretax return is equal to the cost of debt therefore both EPS are equal
pretax return is higher than cost of debt therefore Lenow EPS is lower than Hall EPS
Therefore EPS is unaffected when cost of debt = EBIT/TA =8%
Answer c
Break even level if debt cost increased to 10%
At break even (with 10% cost of debt) we should have EBIT/TA = 10%
EBIT = 10%* TA
EBIT = 10%*900000
EBIT = 90000
Breakeven EBIT = 90000